# Collision-inclusive Manipulation Planning for Occluded Object Grasping via Compliant Robot Motions

### v1

@@ IV-B2 Optimization-based Path Refinement via Collision Minimization
In the second step, we aim to minimize collisions between the gripper and the environment, while still keeping the fingertip following the constraint-exploiting path found in the previous step. For this, we define a collision cost similar to CHOMP [29]. Given the signed distance $\phi(\bm{p})\in\mathbb{R}$ of a point $\bm{p}\in\mathbb{R}^{3}$, the collision cost is defined as follows:
@@ IV-C Impedance Control Generation
where $\bm{\Lambda}\left(\bm{q}_{t}\right)=\bm{J}\left(\bm{q}_{t}\right)^{-\top}\bm{M}\left(\bm{q}_{t}\right)\bm{J}\left(\bm{q}_{t}\right)^{-1}$ is the effective inertia matrix of robot in the task space. When the robot moves locally with a slow speed, we can ignore the effect of term $\dot{\bm{J}}\left(\bm{q}_{t}\right)\dot{\bm{q}}_{t}$. Then, we can obtain a highly approximated yet computationally less expensive dynamics equation, to heuristically guide the generation of impedance controls:
@@
The external wrench $\bm{f}_{ext}\in\mathbb{R}^{6}$ is estimated based on the virtual penetration along the planned geometric path. Specifically, when the gripper is at pose $\widetilde{\bm{X}}_{t}$ while following the geometric path, a local force $\bm{f}\left(\bm{p}\right)\in\mathbb{R}^{3}$ can be estimated at each point $\bm{p}\in\mathcal{P}^{s}$ sampled on the gripper’s surface. The force $\bm{f}\left(\bm{p}\right)$ has a magnitude proportional to the penetration distance (i.e., the negative of signed distance) from this point to the environment; the direction of $\bm{f}\left(\bm{p}\right)$ is the same as the normal direction of the gripper surface at this point:
@@
[…]
$\ddot{\bm{x}}_{\tau}$ $=\bm{\Lambda}^{-1}\left(\bm{f}_{\tau}-\bm{f}_{ext}\left(\widetilde{\bm{X}}_{t}\right)\right)$ (14c)
$\bm{f}_{\tau}$ $=\bm{K}\left(\bm{\xi}_{t}-\bm{x}_{\tau}\right)-\bm{D}\dot{\bm{x}}_{\tau}$ (14d)
@@
Finally, by incorporating the approximated dynamics $\Pi$ as a heuristic constraint in Eq. (15d), the sequence of impedance controls can be generated by solving a constrained optimization program as expressed in Eq. (15). As visualized in Fig. 5, the program aims to have the controlled robot gripper follow the constraint-exploiting geometric path $\left\{\widetilde{\bm{X}}_{1},\cdots,\widetilde{\bm{X}}_{T}\right\}$ planned in Sec. IV-B. As such, the program’s objective is to minimize the distance between the planned geometric path and the predicted path from impedance-controlled dynamics.
@@ V Experiments
We conducted experiments on Franka Emika robots with both single-arm and dual-arm setups, to evaluate the effectiveness of our proposed framework against real-world challenges. We selected $9$ objects for evaluation, as shown in Fig. 6. The selected objects greatly differ in their geometric and physical properties such as size, weight, stiffness, etc. We used a model-based 6D pose estimator FoundationPose [30] with CNOS [31] segmenting the objects on the first frame. The visual models were run on a single RTX 3060 GPU with 12 GB memory. Our collision-inclusive planner was implemented in Python with a single thread on a 3.4 GHz AMD Ryzen 9 5950X CPU; the involved optimization programs (i.e., in Sec. IV-B2 and IV-C) were solved using the CasADi framework [32] with IPOPT solver (an interior point method). The length of the planned path or trajectory (i.e., the number of execution steps) was always set to $T=20$.
@@ V-A Grasping Single Object on a Support Surface
With both single-arm and dual-arm setups, we evaluated our proposed framework in grasping a single object on a flat table, as exemplified in Fig. 7. This task requires the robot to successfully insert one of its fingers into the narrow gap between the object and the table plane, where collisions are inevitable. To robustly find the gap and place the finger there, our framework automatically exploited environmental constraints by contacting the environmental objects, as showcased in Fig. 7. Our framework assumes that the environment configuration cannot be greatly changed by the robot, to keep the task configuration consistent as modeled by our framework. Therefore, for the single-arm setup, we constructed an aluminum frame and statically attached it to the table to constrain the motion range of objects; for the dual-arm setup, we commanded one robot arm to block the object from the opposite side of the grasping direction. The dual-arm experiments show promising feasibility of applying our framework in more realistic setups where the desired constraints are not always naturally available and need to be constructed by the robot itself. We ran $10$ trials for each selected object; for each trial, the maximum number of repetitions was set to $K_{\max}=10$. The trial was logged as a failure if the object was not grasped within $10$ repetitions.
@@
Figure 7: Grasping a single object on a flat table surface by our framework, with both single-arm (left) and dual-arm (right) setups. The green lines show the geometric paths planned by the algorithms in Sec. IV-B, to exploit environmental constraints. Guided by the geometric paths and compensating against estimated external forces, desired impedance controls (red trajectories) are generated to compliantly move the gripper. The table shows evaluation results on $9$ selected objects, where “Reps” is the average number of repetitions before success; the solve times of “$A^{*}$”, “Refine”, and “Impd” correspond to the algorithmic steps described in Sec. IV-B1,  IV-B2, and  IV-C respectively.
We reported the experiment results in the table of Fig. 7. In all experiment trials, our framework succeeded with a small number of needed repetitions ($1.4$ for the single-arm and $1.6$ for the dual-arm setups on average), showing the high effectiveness of our framework in such task setups. Grasping the cracker box and the bowl generally required more repetitions than other objects. This is because their thicknesses are roughly equal to the width between robot fingers when wide open, and thus need the gripper to more precisely orient itself for insertion. Furthermore, the cracker box was empty and more easily deformed, which made the manipulation more difficult by introducing more unmodeled uncertainties. Although the task was not always accomplished for the first time, the task repetitions, as desirably contained in a conceptual manipulation funnel, always produced stable outcomes and would eventually reach the grasp. In terms of solve times, the replanning of each repetition required less than $5$ seconds. The $A^{*}$-based positional path-finding (corresponding to Sec. IV-B1) usually ran fast (i.e., less than $0.5$ seconds); however, due to perception errors, the object may be estimated to penetrate the table with a large negative clearance between them. Such perceptual inaccuracy could require our algorithm to run more iterations by increasing $\delta_{g}$ and to more exhaustively search over the entire workspace to find a feasible path, making the runtime of the iterative $A^{*}$ sometimes longer (e.g., more than $6$ seconds).
@@ V-B Grasping from Tightly Stacked Multiple Objects
For this evaluation, we further challenged our framework with more complicated environment configurations, where multiple objects were tightly packed in a limited space. Specifically, rather than inserting one finger beneath the object in Sec. V-A, this setup requires the robot to insert both fingers into separate gaps at the same time, requiring our framework to deal with more complicated collisions with the environment. We constructed three test scenes of grasping a target book with other books (or objects) snugly next to it to occlude the desired grasp configuration, as shown in Fig. 8: 1. From the tray with a single arm; 2. From the shelf with a single arm; 3. Stacked flat on the table with dual arms (the other arm blocking objects from the other side). For each scenario, we ran $10$ trials with a maximum number of repetitions $K_{\max}=10$. To enable the insertion of both fingers, before execution of each trial, the robot was asked to configure its finger so that the width between fingers was similar to the thickness of the target book.
@@ VI Conclusion
In this work, we proposed a collision-inclusive manipulation planning framework for collision-rich tasks using compliant robot motions, where the desired task configuration (e.g., a stable grasp for the target object) is initially occluded. With real-world experiments on various test scenarios, we show that our framework can effectively address practical occluded grasping problems under uncertainties by exploiting environmental constraints and task repetitions. In the future, we plan to upgrade our framework by integrating real-time force-torque feedback and closing the loop more tightly with an MPC-based strategy, to generate more reactive actions.

### v2

@@ IV-B2 Optimization-based Path Refinement via Collision Minimization
In the second step, we aim to minimize collisions between the gripper and the environment, while still keeping the fingertip following the constraint-exploiting path found in the previous step. To this end, in Eq. (6), given the signed distance $\phi(\bm{p})\in\mathbb{R}$ of a point $\bm{p}\in\mathbb{R}^{3}$, we define a collision cost as in CHOMP [31]. The margin parameter $\varepsilon$, set relative to object’s size, helps prevent the gripper from moving too close to obstacles; a larger value of $\varepsilon$ is used when more conservative collision avoidance is encouraged, such as in the presence of significant perception errors.
@@ IV-C Impedance Control Generation
where $\bm{\Lambda}\left(\bm{q}_{t}\right)=\bm{J}\left(\bm{q}_{t}\right)^{-\top}\bm{M}\left(\bm{q}_{t}\right)\bm{J}\left(\bm{q}_{t}\right)^{-1}$ is the effective inertia matrix of robot in the task space. When the robot moves slowly, the term $\dot{\bm{J}}\left(\bm{q}_{t}\right)\dot{\bm{q}}_{t}$ is neglected, yielding a computationally efficient dynamics approximation:
@@
This approximation may be problematic for tasks involving dynamic, high-speed motions or rapid joint accelerations (e.g., near singularities). However, as we always slowly move the robot in contact-rich actions, our assumption of slow motion always holds in relevant tasks.
The external wrench $\bm{f}_{ext}\in\mathbb{R}^{6}$ is estimated based on the virtual penetration along the planned geometric path. The estimate of $\bm{f}_{ext}$ is rough (and does not need to be accurate given other perception and modeling errors), yet remains informative for heuristically guiding the control generation. Specifically, when the gripper is at pose $\widetilde{\bm{X}}_{t}$ while following the geometric path, a local force $\bm{f}\left(\bm{p}\right)\in\mathbb{R}^{3}$ can be estimated at each point $\bm{p}\in\mathcal{P}^{s}$ sampled on the gripper’s surface. The force $\bm{f}\left(\bm{p}\right)$ has a magnitude proportional to the penetration distance (i.e., the negative of signed distance) from this point to the environment; the direction of $\bm{f}\left(\bm{p}\right)$ is the same as the normal direction of the gripper surface at this point:
@@
$\bm{x}_{\tau}$ $=\bm{x}_{t}+\int_{t}^{\tau}\dot{\bm{x}}_{s}\,ds,\quad\dot{\bm{x}}_{\tau}=\dot{\bm{x}}_{t}+\int_{t}^{\tau}\ddot{\bm{x}}_{s}\,ds$ (14a)
$\ddot{\bm{x}}_{\tau}$ $=\bm{\Lambda}^{-1}\left(\bm{f}_{\tau}-\bm{f}_{ext}\left(\widetilde{\bm{X}}_{t}\right)\right)$ (14b)
$\bm{f}_{\tau}$ $=\bm{K}\left(\bm{\xi}_{t}-\bm{x}_{\tau}\right)-\bm{D}\dot{\bm{x}}_{\tau}$ (14c)
@@
Finally, by incorporating the approximated dynamics $\Pi$ as a heuristic constraint in Eq. (15d), the sequence of impedance controls can be generated by solving a constrained optimization program as expressed in Eq. (15). As visualized in Fig. 5, the program aims to have the controlled robot gripper follow the constraint-exploiting geometric path $\left\{\widetilde{\bm{X}}_{1},\cdots,\widetilde{\bm{X}}_{T}\right\}$ planned in Sec. IV-B. As such, the objective is to minimize $SE(3)$ distance between the planned geometric path and the predicted path from impedance-controlled dynamics.
@@ V Experiments
We conducted real-world experiments on Franka Emika robots with both single-arm and dual-arm setups, to evaluate our framework. As shown in Fig. 6, we selected $9$ objects for evaluation, which greatly differ in geometric and physical properties such as size, weight, stiffness, etc. We used a model-based 6D pose estimator FoundationPose [32] with CNOS [33] segmenting the objects in the first frame. The visual models ran a single RTX 3060 GPU (12 GB). Our collision-inclusive planner was implemented in Python with a single thread on a 3.4 GHz AMD Ryzen 9 5950X CPU; the optimization programs (i.e., in Sec. IV-B2 and IV-C) were solved using the CasADi framework [34] with IPOPT solver (an interior point method). The length of path or trajectory was set to $T=20$. The SDF resolution was $\Delta=1cm$, with $\varepsilon=1cm$ in Eq. (6). Alg. 2 started with $\delta_{g}=0mm$ and incremented it by $d\delta_{g}=1mm$ after each iteration.
@@ V-A Grasping Single Object on a Support Surface
With both single-arm and dual-arm setups, we evaluated our proposed framework in grasping a single object on a flat table, as exemplified in Fig. 7. This task requires the robot to successfully insert one of its fingers into the narrow gap between the object and the table plane, where collisions are inevitable. To robustly find the gap and place the finger there, our framework automatically exploited environmental constraints by contacting the environmental objects, as showcased in Fig. 7. Our framework assumes that the environment configuration cannot be greatly changed by the robot. To this end, for the single-arm setup, we constructed an aluminum frame and fixed it to the table to constrain the motion range of objects; for the dual-arm setup, we commanded one robot arm to block the object from the opposite side of the grasping direction. The dual-arm experiments show promising feasibility of applying our framework in more realistic setups where desired constraints are not naturally available and must be created by the robot itself. We conducted $10$ trials per object, with a maximum of $K_{\max}=10$ repetitions per trial. A trial was considered a failure if the object was not grasped within $10$ repetitions.
@@
Figure 7: Grasping a single object on a flat table surface by our framework, with both single-arm (left) and dual-arm (right) setups. The green lines show the geometric paths planned by the algorithms in Sec. IV-B, and the red trajectories are generated impedance controls. The table shows evaluation results on 99 selected objects, where “Reps” is the average number of repetitions before success; the solve times of $A^{*}$”, “Refine”, and “Impd” correspond to the algorithmic steps described in Sec. IV-B1,  IV-B2, and  IV-C respectively.
We reported the experiment results in the table of Fig. 7. In all experiment trials, our framework succeeded with a small number of needed repetitions ($1.4$ for the single-arm and $1.6$ for the dual-arm setups on average), showing the high effectiveness of our framework in such task setups. Grasping the cracker box and the bowl generally required more repetitions than other objects. This is primarily because their thicknesses are close to the gripper’s maximum opening width, requiring more precise alignment of the gripper’s orientation for successful insertion. Furthermore, the cracker box was empty and prone to deformation, while the bowl had a narrow rim not captured in the mesh model, introducing unmodeled contact interactions that increased manipulation difficulty. Although the task was not always accomplished for the first time, the task repetitions, as desirably contained in a conceptual manipulation funnel, always produced stable outcomes and would eventually reach the grasp. In terms of solve times, the replanning of each repetition required less than $5$ seconds. The $A^{*}$-based positional path-finding (corresponding to Sec. IV-B1) usually ran fast (i.e., less than $0.5$ seconds); however, due to perception errors, the object may be estimated to penetrate the table with a large negative clearance between them. Such perceptual inaccuracy could require our algorithm to run more iterations by increasing $\delta_{g}$ and to more exhaustively search over the entire workspace to find a feasible path, making the runtime of the iterative $A^{*}$ sometimes longer (e.g., more than $6$ seconds).
@@ V-B Grasping from Tightly Stacked Multiple Objects
For this evaluation, we further challenged our framework with more complicated environment configurations, where multiple objects were tightly packed in a limited space. Unlike the single-finger insertion in Sec. V-A, this setup required simultaneous insertion of both fingers into separate gaps, introducing more complex collisions with the environment. We constructed three test scenes of grasping a target book with other books (or objects) snugly next to it to occlude the desired grasp configuration, as shown in Fig. 8: 1. From the tray with a single arm; 2. From the shelf with a single arm; 3. Stacked flat on the table with dual arms (the other arm blocking objects from the other side). For each scenario, we ran $10$ trials with a maximum number of repetitions $K_{\max}=10$. To enable the insertion of both fingers, before execution of each trial, the robot was asked to configure its finger so that the width between fingers was similar to the thickness of the target book.
@@
V-C Comparative Evaluation against Selected Baseline
We selected a contact-rich baseline [37] as it aligns more closely with our task objective, i.e., guiding the robot gripper towards a desired configuration through contacts with approximately static environment (and target object). In contrast, other methods (e.g.,  [22, 23, 24]) focus on manipulating the target object by actively reconfiguring its pose, and therefore may be unsuitable in highly confined spaces where object motion is restricted or undesired.
Refer to captionCase 1Case 2Case 3
Case 1 Case 2 Case 3
Success Time (sec) Success Time (sec) Success Time (sec)
0 / 5 27.99 2 / 5 29.61 0 / 5 32.95
Figure 9: Failure executions by the baseline method.
We adapted the baseline implementation to our task setup using the same approximated dynamics model as in Eq. (14). We found the baseline more sensitive to modeling inaccuracies, making it less reliable when integrated with the highly approximated dynamics. To avoid potential damage to objects or the other robot arm, we limited evaluation to three simple yet representative test cases, with $5$ trials each. Each trial was given a $120$-second time budget, during which the baseline could replan and search for improved solutions, allowing multiple repetitions if time remains. As showcased in Fig. 9, the baseline was more prone to execution failures, such as unsuccessful finger insertion when grasping the book (Case 1 and 3), or incorrect grasps for the clamp (Case 2). Additionally, the baseline’s random sampling-based motion generation required searching in a larger problem space and was inefficient, typically requiring over $20$ seconds to find a solution, as reported in the table in Fig. 9.
@@ VI Conclusion
In this work, we proposed a collision-inclusive manipulation planning framework for collision-rich tasks using compliant robot motions, where the desired task configuration (e.g., a stable grasp for the target object) is initially occluded. With real-world experiments on various test scenarios, we show that our framework can effectively address practical occluded grasping problems under uncertainties by exploiting environmental constraints and task repetitions.